OpenStudy (anonymous):

A ball having a mass of 500 grams is dropped from a height of 9.00 meters. What is its kinetic energy when it hits the ground? A. 4.4 x 10^3 B. 441 C. 44.1 D 4.41

OpenStudy (sleepyhead314):

Hi @jms4323 :) It's been a while since I've tried doing a Physics problem but I think the Equation you need is the one that Converts Potential Energy into Kinetic Energy m*g*h = (1/2)*m*v^2 because the Ball starts at a certain height above the ground which means that it has Potential Energy In this case, the Kinetic Energy would be equal to the Potential Energy because no energy should be lost does that make sense so far? :)

OpenStudy (sleepyhead314):

this means that KE = mgh where m = mass g = gravity (acceleration) h = height

OpenStudy (sleepyhead314):

but be careful! the units for each should be mass = kilograms gravity = meters per seconds squared height = meters as we can see, the Mass that they gave us is Not in the kilograms that we need!

OpenStudy (sleepyhead314):

can you convert 500 grams to kilograms? what would you get?

OpenStudy (sleepyhead314):

once you get that, all you would have to do is plug that into the equation PE = KE = m*g*h where m = 0.5 g = 9.81 h = 9 so 0.5 * 9.81 * 9 = [answer]

OpenStudy (sleepyhead314):

feel free to tag me if you get confused at any part ^_^

OpenStudy (anonymous):

i got the answer much appreciated!!